# Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers

### v1

@@ III.2 Prediction Accuracy: NMR Chemical Shifts
In our study, we utilized the QM9NMR dataset, which contains approximately 134K small organic molecules containing C, N, O, and F (excluding H), with each molecule having no more than nine atoms.[29, 67] This dataset provides the detailed NMR chemical shielding constants for these molecules. To analyze how the model accuracy changes with training data size, we adopted an approach similar to that used in the original publication of the QM9NMR dataset.[29] Specifically, for C, of a total of 831K data points, we randomly withheld 50K data points to build our test set. Subsequently, from the remaining C NMR chemical shifts, we randomly selected subsets containing 100, 200, 500, 1K, 2K, 5K, 10K, 50K, 100K, and 200K data points to create various training sets. For the other isotopes (i.e., H, N, O, and F), the test sets were similarly established by withholding 50K, 30K, 50K, and 1K data points, respectively. The training size for F was set to 2K, whereas the other isotopes were trained on datasets of 100K data points. In addition to the QM9 NMR dataset, we sought to validate the performance of our model on external datasets. Hence, we employed the two sets of molecules provided in another study;[29] one consisting of 40 drug molecules from the GDB17 universe and another containing 12 drugs with 17 or more heavy atoms.
@@
Figure 2 shows the relationship between the mean absolute error (MAE) for C NMR shielding constant predictions and the training data size. Both QKRR and KRR demonstrated consistent improvements in predictive accuracy with an increase in training size. Notably, the quantum kernel exhibited a performance comparable to that of the Laplacian kernel.
For a training size of 100K, the MAE for the 13C predictions was . In a comparative study by Gupta et al., the KRR models using the Coulomb matrix (CM),[89] SOAP, and FCHL descriptors reported MAEs of approximately 4, 2.1, and , respectively, for the same training size.[29]
Compared with the CM descriptor, our GNN-TF descriptor showed significantly better predictive capabilities, achieving an MAE that was nearly half that of the CM descriptor. Although our method did not exceed the accuracy levels of SOAP and FCHL, the performance of the GNN-TF descriptor was competitive, highlighting its potential as a robust descriptor.
| GNN-TF descriptor/Regressor | RMSE (ppm) |
| SchNet/NN222 The value is taken from [90] | 12.8 |
| MEGNet/KRR | 20.4 |
| M3GNet/KRR | 10.0 |
Next, we compared the performance of the GNN-TF descriptors derived from different IAP architectures.
Recently, independent of our work, a predictive model for C NMR chemical shielding was proposed using a pretrained IAP known as SchNet, which is a pioneering GNN used as a descriptor.[90] This model was trained on 400 data points of C NMR chemical shielding constants of the molecules in QM9 dataset,[67] with the SchNet GNN-TF descriptor as an input to a feed-forward NN for regression. The predictive accuracy of the SchNet/NN was .
In pursuit of a fair comparison with their model, we applied Kernel Ridge Regression (KRR) using pre-trained M3GNet and MEGNet GNN-TF descriptors, setting our training data size to 400 data points of C NMR chemical shielding constants. The results of this comparative study are summarized in Table II.
The highest predictive accuracy was achieved by the M3GNet/KRR model, with an root mean-squared error of . This value surpassed the achieved by the MEGNet/KRR model and the achieved by the SchNet/NN model. This clearly underscores the superior performance of the M3GNet GNN descriptor over the other GNN-TF descriptors.
These results suggest that the architecture of the GNN-based IAP influences the performance of the GNN-TF descriptor. A more detailed discussion of the nuances of these architectural differences is presented in Section IV.1.
@@
The accuracy of KRR models incorporating the GNN-TF descriptor for NMR chemical shifts was evaluated for molecules containing five elements that were not present in the training dataset. Table III lists the statistical performance metrics for predicting NMR chemical shifts. Across all elements, the MAE for the test set remained below . The MAE for H and F were notably low at and , respectively, indicating a high degree of prediction accuracy for these nuclei in the unseen molecules. The MAE for O, although higher at , still reflects a reasonable predictive capability, given the complexity of the oxygen chemical shifts. The standard deviation (STD) and interquartile range (IQR) values in the table represent the distribution of chemical shifts within the training data, rather than the accuracy of the model itself. Thus, the higher STD and IQR values for O do not indicate a lack of model precision but rather the natural variability inherent in the O chemical shifts within the training data. The MAE/STD ratio can still offer insights into model performance relative to data variability. For example, the relatively low ratio of O (2.21%) suggests that the model predictions are consistent with the diversity of the training data. Similarly, the higher ratios for H and F indicate that the accuracy of the model is commendable, particularly when considering the broad range of chemical shifts represented in the training dataset.
@@
Subsequently, these models were employed to predict the NMR chemical shifts of molecules containing five elements that were not included in the training data. The results are shown in Fig. 3. The MAE for each nucleus were found to be for H, for C, for N, for O, and for F. The remarkably low MAE for H and C underscores the high accuracy of our model for these nuclei, with predictions that closely mirror the calculated values. The model performed well for the more challenging N and O nuclei, where the chemical shifts can be significantly affected by subtle changes in the molecular structure and environment, as indicated by the MAE values. The F nucleus, while having a higher MAE, showed excellent agreement with the DFT/GIAO calculations, suggesting that the model predictions were robust, even for nuclei with typically higher chemical shift ranges. These results demonstrate the strong predictive power and potential of the model as a reliable tool for accurately predicting NMR chemical shifts across a variety of nuclei, even in molecules beyond the scope of the training data.
| H | C | N | O | F | |
@@
| MAE/STD () | 8.93 | 4.38 | 2.86 | 2.21 | 7.78 |
| FCHL333FCHL results are taken from [29]. | M3GNet GNN-TF | |
| 50K QM9 | 1.88 | 2.28 |
| 40 drugs | 3.7 | 3.47 |
| 12 drugs | 4.2 | 4.23 |
We then expanded our assessment to evaluate the predictive ability of our model for molecules larger than those in the QM9 NMR dataset. As such, we incorporate the test sets provided in another study,[29] which comprised 40 drug molecules from the GDB17 universe and another set containing 12 drugs with 17 or more heavy atoms.
Table IV presents the benchmark results for these test sets using our M3GNet GNN-TF descriptor. For comparison, we used the FCHL descriptor from Gupta’s study.[29] To ensure a fair comparison, we employed our M3GNet GNN-TF descriptor model trained on a size of 100K C. For both models, an increased molecular size in the dataset correlated with deterioration of the MAE value. Notably, although our M3GNet GNN-TF descriptor did not match the value achieved by the FCHL descriptor for the QM9 50K test set, our model exhibited an MAE value that was approximately lower for the 40 GDB17 dataset test. The results were nearly identical for the set of 12 drugs with 17 or more heavy atoms, highlighting that the GNN-TF descriptor was less affected by increasing molecular size.
For a detailed comparison, Fig. 4 illustrates the molecule-specific MAE values for both drug test sets. The molecular structures are provided in Appendix B, and more intricate details are available in another study.[29] Our M3GNet GNN-TF descriptor-based prediction model ensured that the highest MAE values for individual molecules across both test sets remained below . Intriguingly, the desflurane molecule, which posed the greatest challenge, showed MAE values of and for the FCHL and M3GNet GNN-TF descriptor models, respectively. This suggests an approximately 80% reduction in the MAE with our descriptor, which is likely attributable to differences in the encompassed descriptor domain.
@@
In summary, the proposed M3GNet GNN-TF descriptor demonstrates the capability of predicting C NMR chemical shifts for molecules outside the training dataset with an accuracy comparable to that of the state-of-the-art FCHL descriptor.
IV Discussion
IV.1 Influence of Architectural Choices on GNN-TF Descriptor Performance
In our exploration of different architectures for generating GNN-TF descriptors, we observed several patterns. SchNet, which operates on GNN-based local descriptors to evaluate systems as summations of atomic energies, accounts only for pairwise interactions. This limited inclusion could potentially constrain expressions, leading to inadequate representational power.
The subpar performance of MEGNet during transfer learning may be attributed to its architectural design as it integrates atomic (local) descriptors into molecular (global) descriptors through concatenation. This means that the final piece of information passed to the MLP is not extracted directly from the end of the model, which might not be the optimal representation for targeted atomic-wise property prediction; however, it is expected to be suitable for molecule-wise property predictions.
IV.2 Significance of Dataset Size and Diversity
@@
The M3GNet training regimen incorporates data from 187,687 ionic steps spanning 62,783 compounds, including 187,687 energies, 16,875,138 force components, and 1,689,183 stress components. This diverse dataset covers 89 elements from the periodic table. The model is not limited to learning only the energies associated with these elements but extends to atomic-level forces. Moreover, M3GNet training includes not only stable structures but also the processes of structure optimization. The ingestion of vast amounts of data from crystalline systems may have endowed the M3GNet with enhanced expression, potentially making it adept at interpolating molecular systems.
IV.3 Potential for Transfer Learning on Quantum Computer
There is a potential for leveraging quantum computation approaches.[91] Specifically, our 10 qubit QKRR, facilitated by a simulator, demonstrated a performance comparable to that of state-of-the-art KRR. This is underpinned by the theoretical equivalence of the NPQC with the RBF.
@@ IV.3 Potential for Transfer Learning on Quantum Computer
The quantum kernel method stands out because of its capability to compute with fewer measurement iterations than other quantum computation methodologies, such as quantum neural networks.[92] In particular, our proposed M3GNet GNN-TF descriptor can be feasibly realized with a minimum of six qubits, enabling evaluations with a quantum bit count that is more efficient than traditional descriptors, such as SOAP. However, embedding for higher-dimensional SOAP appears to be a challenge, possibly due to noise.
From a futuristic perspective, there is excitement about the possibility of developing kernels that traditional computers cannot express, as well as accelerating the inversion calculations of kernel matrices using quantum algorithms. The constant scaling property of our proposed method concerning element number dimensions may significantly contribute to real-time material exploration powered by quantum-classical hybrid algorithms in the near future.
V Conclusion
@@
The dynamics of machine learning and its extensive applications across various domains are driving cutting-edge research. Our endeavor to integrate transfer learning with pretrained IAP GNNs for NMR chemical shift prediction offers a paradigm shift in efficiency and scalability. The GNN-TF descriptor presents an unparalleled advantage in terms of scalability due to its consistent dimensionality, irrespective of the number of elements.
@@
Comparative evaluations with other renowned descriptors, such as SOAP, suggest that the GNN-TF descriptor can match, if not surpass, the performance of its contemporaries while maintaining a more compact representation. This is especially important when factoring large datasets, where dimensionality can exponentially burgeon.
@@
Architectural choice plays a pivotal role in the performance of GNN-TF descriptors. Moreover, the diversity and vastness of the training dataset, which encompasses myriad elemental types and structural configurations, augment the robustness and versatility of the GNN.
@@
Note added - As we were finalizing this manuscript, we became aware of recent articles [90, 93, 94] that also utilize intermediate information from graph neural network potentials. In Section III.2, we added a direct comparison between our results and theirs. Elijošius et al. applied the pre-trained MACE descriptor to generative modeling of molecules [94].
Acknowledgements.

### v2

@@ 3.2 Prediction Accuracy: NMR Chemical Shifts
In our study, we utilized the QM9NMR dataset, which contains approximately 134K small organic molecules containing C, N, O, and F (excluding H), with each molecule having no more than nine atoms.29, 82 This dataset provides the detailed NMR chemical shielding constants for these molecules. To analyze how the model accuracy changes with training data size, we adopted an approach similar to that used in the original publication of the QM9NMR dataset.29 Specifically, for 13C, of a total of 831K data points, we randomly withheld 50K data points to build our test set. Subsequently, from the remaining 13C NMR chemical shifts, we randomly selected subsets containing 100, 200, 500, 1K, 2K, 5K, 10K, 50K, 100K, and 200K data points to create various training sets. For the other isotopes (, 1H, 15N, 17O, and 19F), the test sets were similarly established by withholding 50K, 30K, 50K, and 1K data points, respectively. The training size for 19F was set to 2K, whereas the other isotopes were trained on datasets of 100K data points. In addition to the QM9 NMR dataset, we sought to validate the performance of our model on external datasets. Hence, we employed the two sets of molecules provided in another study;29 one consisting of 40 drug molecules from the GDB17 universe and another containing 12 drugs with 17 or more heavy atoms.
@@
Fig. 2 shows the relationship between the mean absolute error (MAE) for the 13C NMR shielding constant predictions and the training data size. Both QKRR and KRR demonstrated consistent improvements in predictive accuracy with an increase in training size. Notably, the quantum kernel exhibited a performance comparable to that of the Laplacian kernel. For a training size of 100K, the MAE for the 13C predictions was . In a comparative study by Gupta et al., the KRR models using the Coulomb matrix (CM),109 SOAP, and FCHL descriptors reported MAEs of approximately 4, 2.1, and , respectively, for the same training size.29 Compared with the CM descriptor, our GNN-TL descriptor showed significantly better predictive capabilities, achieving an MAE that was nearly half that of the CM descriptor. Although our method did not exceed the accuracy levels of SOAP and FCHL, the performance of the GNN-TL descriptor was competitive, highlighting its potential as a robust descriptor.
| GNN-TL descriptor/Regressor | RMSE (ppm) |
| SchNet/NN333 | 12.8 |
| MEGNet/KRR | 20.08±0.55 |
| M3GNet/KRR | 10.02±0.37 |
| MACE-MP-0-small/KRR | 9.77±0.34 |
| MACE-MP-0-large/KRR | 9.74±0.27 |
| MACE-OFF23-small/KRR | 8.05±0.19 |
| MACE-OFF23-large/KRR | 8.15±0.42 |
Next, we compared the performance of the GNN-TL descriptors derived from different IAP architectures. Recently, independent of our work, a predictive model for 13C NMR chemical shielding was proposed using a pretrained IAP known as SchNet, which is a pioneering GNN used as a descriptor.110 This model was trained on 400 data points of 13C NMR chemical shielding constants of the molecules in QM9 dataset,82 with the SchNet GNN-TL descriptor as an input to a feed-forward NN for regression. The predictive accuracy of the SchNet/NN was a root mean-squared error (RMSE) of . In pursuit of a fair comparison with their model, we applied KRR using pre-trained MEGNet, M3GNet and MACE GNN-TL descriptors, setting our training data size to 400 data points of 13C NMR chemical shielding constants. To account for the influence of random sampling, we created 10 different training sets, each comprising 400 data points. The effect of potential data bias was then quantified by calculating the mean RMSE and standard deviation (STD) for each model. Detailed verification including kernel function dependencies can be found in the Appendix. The results of this comparative study are summarized in Table 2. In Table 2, the results for KRR using the Gaussian kernel, which showed superior accuracy compared to the Laplacian kernel, are presented.
In contrast to the SchNet/NN model’s RMSE of 12.8 ppm, the MEGNet/KRR model shows significantly lower predictive accuracy with an RMSE of 20.08 ± 0.55 ppm, suggesting that the MEGNet descriptor is less effective for 13C NMR chemical shielding data. The M3GNet/KRR model demonstrates a substantial improvement with an RMSE of 10.02 ± 0.37 ppm. Models using MACE descriptors show even greater accuracy: the MACE-MP-0-small/KRR and MACE-MP-0-large/KRR models achieve RMSEs of 9.77 ± 0.34 ppm and 9.74 ± 0.27 ppm, respectively. The best performance is observed with the MACE-OFF23-small/KRR model, which has an RMSE of 8.05 ± 0.19 ppm, with the MACE-OFF23-large/KRR model close behind at 8.15 ± 0.42 ppm. These results highlight the superior performance of the MACE descriptors, particularly MACE-OFF23-small, in enhancing the accuracy of KRR models for predicting 13C NMR chemical shielding. A more detailed discussion of the nuances of these architectural differences is presented in Section 4.1.
@@
The accuracy of KRR models incorporating the M3GNet GNN-TL descriptor with a Laplacian kernel for NMR chemical shifts was evaluated for each test set of the five different nuclei. Table 3 lists the statistical performance metrics for predicting NMR chemical shifts. Across all elements, the MAE for the test set remained below . The MAE for 1H and 19F were notably low at and , respectively, indicating a high degree of prediction accuracy for these nuclei in the unseen molecular environments. The MAE for 17O, although higher at , still reflects a reasonable predictive capability, given the complexity of the oxygen chemical shifts. The STD and interquartile range (IQR) values in the Table 3 represent the distribution of chemical shifts within the training data, rather than the accuracy of the model itself. Thus, the higher STD and IQR values for 17O do not indicate a lack of model precision but rather the natural variability inherent in the 17O chemical shifts within the training data. The MAE/STD ratio can still offer insights into model performance relative to data variability. For example, the relatively low ratio of 17O (2.21%) suggests that the model predictions are consistent with the diversity of the training data. On the other hand, the higher ratios for 1H (9.09%) and 19F (7.78%) indicate that the accuracy of the models are not as high as desired, particularly when considering the range of chemical shifts represented in the training dataset. The maximum absolute error (MaxAE) for all nuclei is comparable to the STD of the training data. This is attributed to random sampling and is expected to improve with the application of more sophisticated data point sampling techniques, such as active learning.
@@
Subsequently, these models were employed to predict the NMR chemical shifts of a single molecule containing five elements that was not included in the training data. The results are shown in Fig. 3. The MAE for each nucleus were found to be for 1H, for 13C, for 15N, for 17O, and for 19F. The remarkably low MAE for 1H and 13C underscores the high accuracy of our model for these nuclei, with predictions that closely mirror the calculated values. The model performed well for the more challenging 15N and 17O nuclei, where the chemical shifts can be significantly affected by subtle changes in the molecular structure and environment, as indicated by the MAE values. The 19F nucleus, while having a higher MAE, showed excellent agreement with the DFT/GIAO calculations, suggesting that the model predictions were robust, even for nuclei with typically higher chemical shift ranges. These results demonstrate the strong predictive power and potential of the model as a reliable tool for accurately predicting NMR chemical shifts across a variety of nuclei, even in molecules beyond the scope of the training data.
| 1H | 13C | 15N | 17O | 19F | |
@@
| MAE/STD () | 9.09 | 4.38 | 2.86 | 2.21 | 7.78 |
| FCHL444 |
| 50K QM9 | 1.88 | 2.28 (68.58) | 1.87 (59.76) | ||||
| 40 drugs | 3.7 | 3.46 (29.86) | 2.83 (16.08) | ||||
| 12 drugs | 4.2 | 4.21 (20.48) | 3.85 (24.70) |
We then expanded our assessment to evaluate the predictive ability of our model for molecules larger than those in the QM9 NMR dataset. As such, we incorporate the test sets provided in Ref. 29, which comprised 40 drug molecules from the GDB17 universe and another set containing 12 drugs with 17 or more heavy atoms. See Ref. 29 for the structures of these molecules.
Table 4 presents the benchmark results for each test set using our M3GNet GNN-TL descriptor and MACE-OFF23-small GNN-TL descriptor. For comparison, we used the FCHL descriptor from Gupta’s study.29 To ensure a fair comparison, we employed our GNN-TL descriptor models trained on a size of 100K 13C chemical shielding constants. For both models, an increased molecular size in the dataset correlated with deterioration of the MAE value. Notably, although our M3GNet GNN-TL descriptor did not match the value achieved by the FCHL descriptor for the QM9 50K test set, our model exhibited an MAE value that was approximately lower for the 40 GDB17 dataset test. The MACE-OFF23-small GNN-TL descriptor showed even better performance, with an MAE of for the QM9 50K test set, closely matching the FCHL descriptor, and significantly outperforming it for the 40 GDB17 dataset with an MAE of . For the set of 12 drugs with 17 or more heavy atoms, the M3GNet descriptor showed an MAE of , while the MACE-OFF23-small descriptor showed an MAE of . Notably, the M3GNet descriptor’s accuracy is comparable to the FCHL descriptor. The results were nearly identical for the set of 12 drugs with 17 or more heavy atoms, highlighting that the M3GNet GNN-TL descriptor was less affected by increasing molecular size. On the other hand, the MACE-OFF23-small descriptor significantly outperforms FCHL with an MAE of , highlighting its superior predictive performance.
For a detailed comparison, Fig. 6 illustrates the molecule-specific MAE values for both drug test sets. The molecular structures are provided in Ref.29. Our M3GNet and MACE-OFF-small GNN-TL descriptor-based prediction models ensured that the highest MAE values for individual molecules across both test sets remained below . Intriguingly, the desflurane molecule, which posed the greatest challenge, showed MAE values of , and for the FCHL, M3GNet and MACE-OFF23-small GNN-TL descriptor models, respectively. This suggests an approximately 80% reduction in the MAE with our descriptor, which is likely attributable to differences in the encompassed descriptor domain.
@@
Lastly, to explore further practical applications of the constructed models, we validated the NMR chemical shielding constants obtained using semi-empirical PM7-level geometries as inputs against the NMR chemical shift values obtained using DFT/GIAO-level structures from the training data. This validation was performed on the QM9 50K holdout set and two drug molecule test sets, as provided by Ref.29. The 13C prediction model employed was the M3GNet/KRR model. The MAE values for each molecule in the drug datasets can be found in Figures 6(b) and 6(d). For the QM9 50K holdout set, the result was , showing a significant deterioration of compared to when DFT-level geometries were used as inputs. Conversely, predictions for the 40 drugs and 12 drugs test sets showed only minor deteriorations of and , respectively. These results suggest that even when using more readily available PM7-level geometries as inputs, the transferability of the model remains robust for extrapolative predictions on larger molecules compared to the training data.
4 Discussion
4.1 Influence of Architectural Choices on GNN-TL Descriptor Performance
In our exploration of different architectures for generating GNN-TL descriptors, we observed several patterns. First, as shown in Table 1 and Table 2, it is important to note that the accuracy of GNN-TL descriptors does not necessarily improve with an increase in the dimensionality of the descriptors. With this in mind, we discuss the architecture of each GNN-based IAP. SchNet, which operates on GNN-based local descriptors to evaluate systems as summations of atomic energies, accounts only for pairwise interactions. This limited inclusion could potentially constrain expressions, leading to inadequate representational power. The subpar performance of MEGNet during transfer learning may be attributed to its architectural design as it integrates atomic (local) descriptors into molecular (global) descriptors through concatenation. This means that the final piece of information passed to the MLP is not extracted directly from the end of the model, which might not be the optimal representation for targeted atomic-wise property prediction; however, it is expected to be suitable for molecule-wise property predictions. Moreover, the M3GNet architecture, which considers three-body interactions, has the potential to capture the three-dimensional structure of molecules with high resolution. Additionally, the MACE model, an E(3) equivariant GNN, has demonstrated high performance as an IAP, suggesting that the outputs of its GNN layers are highly accurate in representing molecular structures. Furthermore, future improvements in accuracy may be achieved by leveraging the outputs of higher-order GNN layers in the MACE model, corresponding to the two-body and three-body terms in the atomic cluster expansion.
4.2 Significance of Dataset Size and Diversity
@@
The M3GNet training regimen incorporates data from 187,687 ionic steps spanning 62,783 compounds, including 187,687 energies, 16,875,138 force components, and 1,689,183 stress components. This diverse dataset covers 89 elements from the periodic table. The model is not limited to learning only the energies associated with these elements but extends to atomic-level forces. Moreover, M3GNet training includes not only stable structures but also the processes of structure optimization. The ingestion of vast amounts of data from crystalline systems may have endowed the M3GNet with enhanced expression, potentially making it adept at interpolating molecular systems. The pre-trained MACE-MP0 model was trained using ten times more energy data of crystalline systems, potentially contributing to the improved accuracy of the 13C NMR chemical shift predictions shown in Table 2. On the other hand, the MACE-OFF23 model, which is specialized for molecules containing 10 elemental species, was trained on a dataset comprising about 1M energy data points, with structures containing up to 150 atoms. This extensive training dataset might make it more suitable for predicting molecular NMR chemical shifts. Thus, the training data for IAPs, much like their architectures, could be a crucial factor in determining the performance of the descriptors.
4.3 Potential for Transfer Learning on Quantum Computer
@@
There is a potential for leveraging quantum computation approaches.111 Specifically, our 10 qubit QKRR, facilitated by a simulator, demonstrated a performance comparable to that of state-of-the-art KRR. This is underpinned by the theoretical equivalence of the NPQC with the Gaussian kernel. The quantum kernel method stands out because of its capability to compute with fewer measurement iterations than other quantum computation methodologies, such as quantum neural networks.112 In particular, our proposed M3GNet GNN-TL descriptor can be feasibly realized with a minimum of six qubits, enabling evaluations with a quantum bit count that is more efficient than traditional descriptors, such as SOAP. However, embedding for higher-dimensional SOAP appears to be a challenge, possibly due to noise. From a futuristic perspective, there is excitement about the possibility of developing kernels that traditional computers cannot express, as well as accelerating the inversion calculations of kernel matrices using quantum algorithms. The constant scaling property of our proposed method concerning element number dimensions may significantly contribute to real-time material exploration powered by quantum-classical hybrid algorithms in the near future.
5 Conclusion
@@
The dynamics of machine learning and its extensive applications across various domains are driving cutting-edge research. Our endeavor to integrate transfer learning with pretrained IAP GNNs for NMR chemical shift prediction offers a paradigm shift in efficiency and scalability. The GNN-TL descriptor presents an unparalleled advantage in terms of scalability due to its consistent dimensionality, irrespective of the number of elements.
@@
Comparative evaluations with other renowned descriptors, such as SOAP, suggest that the GNN-TL descriptor can match, if not surpass, the performance of its contemporaries while maintaining a more compact representation. This is especially important when factoring large datasets, where dimensionality can exponentially burgeon.
@@
Architectural choice plays a pivotal role in the performance of GNN-TL descriptors. Moreover, the diversity and vastness of the training dataset, which encompasses myriad elemental types and structural configurations, augment the robustness and versatility of the GNN.
@@
Note added - As we were finalizing this manuscript, we became aware of recent articles 110, 113, 86 that also utilize intermediate information from graph neural network potentials. In Section 3.2, we added a direct comparison between our results and theirs. Elijošius et al. applied the pre-trained MACE descriptor to generative modeling of molecules 86.
Appendix
Distribution of Datasets for Each NMR Chemical Shift Prediction Model
The distributions of the training and test sets sampled from the QM9NMR dataset are shown in Fig. 4. Fig. 4(a) shows that above 5K, the distribution is in good agreement with the overall distribution of the 13C NMR shielding constants. For the other elemental species, the distributions of the training and test sets were in good agreement with the overall distribution.
Kernel Function Dependency for Various GNN-TL Descriptors
| GNN-TL descriptor | Gaussian kernel | Laplacian kernel |
| MEGNet | 20.08±0.55 | 21.12±0.56 |
| M3GNet | 10.02±0.37 | 10.31±0.38 |
| MACE-MP-0-small | 9.77±0.34 | 10.78±0.31 |
| MACE-MP-0-large | 9.74±0.27 | 10.17±0.30 |
| MACE-OFF23-small | 8.05±0.19 | 8.64±0.13 |
| MACE-OFF23-large | 8.15±0.42 | 8.77±0.21 |
The accuracy of KRR models using Gaussian and Laplacian kernels was evaluated. Table 5 presents the mean RMSE and its standard deviation for predictions on the 50K holdout set by models trained on 400 data points of ¹³C using Gaussian and Laplacian kernels. For all models using GNN-TL descriptors, the mean RMSE of models with Gaussian kernel was found to be more accurate than those with Laplacian kernel. However, the variation in accuracy due to dataset sampling (standard deviation) was found to have a greater impact than kernel choice in models with MEGNet and M3GNet GNN-TL descriptors. On the other hand, in models with MACE GNN-TL descriptors, the impact of kernel choice was more significant than the variation due to dataset sampling.
Next, Table 6 shows the accuracy of KRR models using M3GNet and MACE-OFF23-small GNN-TL descriptors trained on a 100K ¹³C training set. Unlike models trained on the 400 ¹³C training set, the KRR models with M3GNet GNN-TL descriptors consistently showed higher accuracy with the Laplacian kernel compared to the Gaussian kernel. Conversely, the results for MACE-OFF23-small GNN-TL descriptors were similar to those for models trained on the 400 ¹³C training set, with the Gaussian kernel models demonstrating higher accuracy. This suggests that the appropriate kernel function may vary depending on the size of the training data.
Finally, these results indicate the choice of kernel functions for KRR models as presented in the Results section of this paper. For models trained on 400 ¹³C data points, all KRR models using GNN-TL descriptors employed the Gaussian kernel. In contrast, for models trained on 100K ¹³C data points, the Laplacian kernel was used for KRR models with M3GNet GNN-TL descriptors, whereas the Gaussian kernel was employed for models with MACE-OFF23-small GNN-TL descriptors.
| M3GNet | MACE-OFF23-small | |||
| Gaussian | Laplacian | Gaussian | Laplacian | |
| 50K QM9 | 2.35 | 2.28 | 1.87 | 2.10 |
| 40 drugs | 3.98 | 3.46 | 2.83 | 3.21 |
| 12 drugs | 5.14 | 4.21 | 3.85 | 3.93 |
Validation of Learning Accuracy of NMR Chemical Shift Prediction
Fig. 5 illustrates the accuracy of the KRR models trained using the M3GNet GNN-TL descriptor for five elemental species. The MAE values for the NMR shielding constant, for train/test, are as follows: for 1H, 0.0344/0.1767; 13C, 0.1420/2.2798 ppm; 15N, 0.3910/3.4157 ppm; 17O 0.8881/4.9509 ppm; and 19F 0.0864/2.6518 ppm.
The accuracy of the GNN-TL descriptors was also validated using the molecular structures of two drug molecule data sets reported in Ref 29. The predicted 13C NMR shielding constants for each drug molecule using the M3GNet and MACE-OFF23 GNN-TL/KRR models are shown in Fig. 6 (a) and (c). These predictions are accompanied by the values predicted by the FCHL/KRR model.29 The prediction results of the M3GNet/KRR model using PM7-level optimized geometries, along with the prediction results using DFT-level geometries, are shown in Fig. 6(b) and (d).
Data Availability
Data and code required to reproduce the figures and tables related to the GNN-TL descriptors and the NMR shielding constants prediction models presented in the manuscript is publicly accessible on GitHub at https://github.com/TShiotaSS/gnn-tl. The dataset utilized for the prediction of NMR chemical shifts, specifically the QM9NMR dataset, is available at DOI: https://doi.org/10.17172/NOMAD/2021.10.16-1 and GitHub at https://moldis-group.github.io/qm9nmr/. Results of the DFT/GIAO calculations for isolated atoms, used for NMR chemical shift computations, are included within the manuscript. The GNN-TL descriptor vectors for the QM9NMR datasets are available at DOI: https://doi.org/10.6084/m9.figshare.25484068.v2. We have modified the code to extract GNN-TL descriptors from the pretrained M3GNet model on Github at https://github.com/materialsvirtuallab/m3gnet, and this adapted version can be found at https://github.com/TShiotaSS/gnn-tl/tree/main/scripts/m3gnet. The code used to extract GNN-TL descriptors from the pretrained MEGNet model can be found on GitHub at https://github.com/materialsvirtuallab/megnet/blob/master/megnet/utils/descriptor.py. The code used to generate descriptors from the pretrained MACE models can be found on GitHub at https://github.com/ACEsuit/mace/blob/main/mace/calculators/mace.py. The implementation for quantum kernel ridge regression used in this study is available at https://github.com/Qulacs-Osaka/scikit-qulacs/tree/main/skqulacs/qkrr.
Conflicts of interest
There are no conflicts to declare.
Acknowledgements